# SHARE: Shared Memory-Aware Open-Domain Long-Term Dialogue Dataset Constructed from Movie Script

### v1

@@ 5.3 Case Study
We examined how shared memory influences interactions between dialogue models. Similar to the previous multi-session evaluation, we analyzed how two models select memory and generate responses when provided with memory and two conversations. The models must choose relevant memory based on the dialogue and generate appropriate responses. Figure 3 illustrates how the proposed EPISODE framework utilizes user information and shared memories between participants.
Without shared memory: In this dialogue, the model generates appropriate responses by reflecting each speaker’s persona. However, it focuses only on the persona and does not include any specific events or shared experiences between the two speakers. As a result, only the basic aspects of their relationship are revealed, and the conversation lacks a strong sense of connection between the speakers due to the absence of references to unique shared events.
With shared memory: In this dialogue, the model similarly generates fitting responses, but with a key difference: The model refers to past events shared by the speakers, such as, “Do you remember when we made that bet about me getting my driver’s license?”. This indicates that the model effectively utilizes shared memory, highlighting shared experiences between the speakers. Using shared memory makes the conversation more engaging and deeper, allowing the relationship and emotions between speakers to emerge naturally, offering the user a more immersive experience.
@@ 6 Conclusion
In this study, we introduce SHARE, a new long-term dialogue dataset that includes shared experiences between two participants. This dataset, extracted from movie scripts, provides a critical foundation for the development of long-term dialogue systems. We propose EPISODE, a novel approach for long-term dialogue systems, which promotes more natural conversations. Our experimental results confirm that our memory update method is more effective than simply accumulating memory. Furthermore, by integrating shared memory into the dialogue system, we demonstrate improvements in both reflectiveness and consistency, leading to richer and more coherent interactions. By making the newly collected SHARE dataset publicly available, we aim to contribute to future research. This study offers a new direction for the development of long-term dialogue systems and is expected to serve as a crucial stepping stone in enhancing conversation quality.
@@
We create a dialogue dataset of two individuals with shared memories based on movie scripts. Despite the strong performance demonstrated by the models trained on SHARE, there are several limitations to consider. First, since the dataset is constructed from movie scripts, the model may generate dramatic dialogues that differ from typical chit-chat. Second, the use of imperfect machine learning models (e.g. GPT-4 and GPT-3.5-turbo) for information extraction and annotation may result in inaccuracies in this information and annotations.
@@
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00350430) and by Smilegate AI Center.
@@ A.1 Dataset Collection
We use GPT-4 to extract information from movie scripts. The detailed prompt is provided in the Table 10. Before annotation, we parsed the Persona sentences because they contained too much information in a single sentence. The prompts for parsing and annotation are provided in Table 11 and Table 12. We used GPT-3.5-turbo for both parsing and annotation.
@@ A.2 Implementation and Training Details
Response Generator We trained the Response Generator model using both the Llama and Gemma models.
We train using the Gemma model on 2 NVIDIA RTX A5000 GPU devices, with a cosine scheduler, starting with a learning rate of $1\times 10^{-5}$ and using 1000 warmup steps. The per-device batch size is set to 2, and gradients are accumulated over 1 step. The maximum output sequence length is 1024 tokens. The model is trained for 4 epochs. It takes 2 hours to train the model.
We train using the Llama 3 Dubey et al. (2024) model on 2 NVIDIA RTX A6000 GPU devices, with a cosine scheduler, starting with a learning rate of $1\times 10^{-5}$ and using 1000 warmup steps. The per-device batch size is set to 8, and gradients are accumulated over 1 step. The maximum output sequence length is 1024 tokens. The model is trained for 4 epochs. It takes 8 hours to train the model.
Information extraction module We use the Llama 3  Dubey et al. (2024) model with 2 NVIDIA RTX A6000 GPUs and a batch size of 4, with a training time of 8 hours. The maximum output sequence length is 2048 tokens.
Memory selection module We use the Llama 3  Dubey et al. (2024) model with 2 NVIDIA RTX A6000 GPUs and a batch size of 4, with a training time of approximately 2 days. The maximum output sequence length is 2048 tokens.
Memory update module We train using the Llama 3 Dubey et al. (2024) model on 2 NVIDIA RTX A6000 GPU devices, with a cosine scheduler, starting with a learning rate of $1\times 10^{-5}$ and using 1000 warmup steps. The per-device batch size is set to 4 for both training and evaluation, and gradients are accumulated over 1 step. The maximum length of the output sequence is 2048 tokens. It takes 5 hours to train the model.
@@ A.3 Memory selection
In memory selection, the most appropriate memory must be chosen to ensure contextually relevant dialogue. To select the appropriate memory, it is necessary to consider all memories from previous sessions and all utterances from the current conversation. To prevent the model from being influenced by the order of the memories, and to mitigate any potential bias caused by sequence ordering, the memories were shuffled before being processed. The prompt used for memory selection is presented in Table 13.
@@ A.4 Information extraction
We conducted experiments to evaluate how similar and consistent the model’s output is compared to the actual ground truth. The experiments were performed using the BLEU-3/4 Papineni et al. (2002), ROUGE (ROUGE-1/2/L) Lin (2004), and BERTscore Zhang et al. (2019) metrics, and the model was compared against GPT-3.5-turbo. The results are presented in Table 6. The prompt used during training is very similar to the one used in the information extraction part of Session 3. However, in the training prompt, the phrase “from the movie {movie_name}. Use your knowledge of the characters to enhance your analysis where applicable.” was removed. The full details of the modified prompt can be found in Table 10.
@@ A.6 Experiments
Evaluating the performance of open-domain conversations is critical for advancing natural language understanding. While traditional metrics remain important, recent methods have incorporated assessments utilizing models like GPT-4 to enhance evaluation accuracy. One such method, GPT Eval, leverages Chain of Thought (CoT) reasoning to evaluate a model’s text generation. This technique offers greater context and structured guidance for large language models (LLMs) when assessing generated text. For this evaluation, we selected multi-session dialogues from the SHARE dataset, each consisting of more than five sessions. From each session, two utterances were extracted to create dialogues within an EPISODE. The experiment was conducted five times for each criterion, and the results were averaged to determine the final outcome. The GPT Eval Liu et al. (2023) prompt used in this process is provide in Table 17, Table 18, Table 19, Table 20, Table21, Table 22, Table 23.
@@ A.7 Human Evaluation
We conduct human evaluations through Amazon Mechanical Turk to highlight the differences between Share Model and Noshare Model. We show the interface for the evaluation in Figure 4. We randomly select 25 conversations from 96 test sets, using a random seed of 2025, and have 20 annotators evaluate each conversation based on the following metrics: closeness, engagingness, and coherence. The evaluation takes place in the third session to maintain consistency in the testing environment. Table 7 provides a comparison between the Share and Noshare conversations, showing that the Share model achieves superior results over the Noshare model. This demonstrates that conversations with Shared Memory are more engaging and result in more natural dialogue.

### v3

@@ 5.3 Case Study
We examine how shared memory influences interactions between dialogue models. Figure 3 presents example dialogues generated by the proposed EPISODE framework, comparing cases with and without shared memory usage.
Without shared memory. In this dialogue, the model generates responses that align with each speaker’s persona. However, it relies solely on persona information and does not incorporate any specific shared events or experiences between the speakers. Consequently, while the conversation maintains coherence, it lacks depth and a strong sense of connection, as it fails to reference unique moments that define their relationship.
With shared memory. In this dialogue, the model integrates shared experiences, exemplified by references such as, "Do you remember when we made that bet about me getting my driver’s license?" This demonstrates the model effectively utilizes shared memory, reinforcing past interactions between the speakers. By incorporating shared experiences, the conversation becomes more engaging and dynamic, naturally revealing the relationship and emotions between the participants. Moreover, the dialogue conveys a sense of familiarity and warmth, enhancing the authenticity of their bond and providing a more immersive user experience.
@@ 6 Conclusion
In this study, we introduce SHARE, a new long-term dialogue dataset that includes shared experiences between two participants. This dataset, extracted from movie scripts, provides a critical foundation for the development of long-term dialogue systems. We propose EPISODE framework, a novel approach for long-term dialogue systems, which promotes more natural conversations. We also utilize a novel evaluation method, GPT-Eval, to extend beyond traditional metrics and incorporate diverse evaluation criteria. By integrating shared memory into the dialogue system, we demonstrate improvements in both reflectiveness and consistency, which lead to richer and more coherent interactions. By making the newly collected SHARE dataset publicly available, we aim to contribute to future research.
@@
We create a dialogue dataset of two individuals with shared memories based on movie scripts. Despite the strong performance demonstrated by the models trained on SHARE, there are several limitations to consider. First, since the dataset is constructed from movie scripts, the model may generate dramatic dialogues that differ from typical chit-chat. Second, the use of imperfect machine learning models (e.g. GPT-4 and GPT-3.5-turbo) for information extraction and annotation may result in inaccuracies in this information and annotations. Third, using multiple GPT models during dataset construction may introduce inconsistencies in the extracted information. Initially, GPT-4 is used for information extraction, but later processes, such as memory updates and utterance mapping, employ GPT-4o and GPT-3.5-turbo to reduce costs.
@@
Acknowledgment
This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP)-ICT Creative Consilience Program grant funded by the Korea government(MSIT)(IITP-2025-RS-2020-II201819) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00350430). Finally, this research was also supported by the Smilegate AI Center.
@@ A.1 Dataset Collection
We use GPT-4 to extract information from movie scripts. The detailed prompt is provided in the Table 11. Before annotation, we parse the Persona sentences because they contained too much information in a single sentence. The prompts for parsing and annotation are provided in Table 12 and Table 13. We use GPT-3.5-turbo for both parsing and annotation. In this study, we divide the dataset sessions into training(14,255), testing(1,685), and validation(1,739) sets. The distribution of dataset genres is shown in the Figure 5
@@ A.2 Implementation and Training Details
Response Generator We train the Response Generator model using both the Llama and Gemma models. We train the Gemma on 2 NVIDIA RTX A5000 GPU devices, with a cosine scheduler, starting with a learning rate of $1\times 10^{-5}$ and using 1000 warmup steps. The per-device batch size is set to 2, and gradients are accumulated over 1 step. The maximum output sequence length is 1024 tokens. The model is trained for 4 epochs. It takes 2 hours to train the model. We train the Llama on 2 NVIDIA RTX A6000 GPU devices, with a cosine scheduler, starting with a learning rate of $1\times 10^{-5}$ and using 1000 warmup steps. The per-device batch size is set to 8, and gradients are accumulated over 1 step. The maximum output sequence length is 1024 tokens. The model is trained for 4 epochs. It takes 8 hours to train the model.
We employ Low-Rank Adaptation (LoRA)Hu et al. (2021), a computationally efficient tuning method, for both the LLAMA and Gemma models. LoRA reduces the trainable parameters by utilizing pairs of rank-decomposition matrices. In our experiments, we set the rank to 16
Information extraction module We use the Llama with 2 NVIDIA RTX A6000 GPUs and a batch size of 4, with a training time of 8 hours. The maximum output sequence length is 2048 tokens.
Memory selection module We use the Llama with 2 NVIDIA RTX A6000 GPUs and a batch size of 4, with a training time of approximately 2 days. The maximum output sequence length is 2048 tokens.
Memory update module We train the Llama on 2 NVIDIA RTX A6000 GPU devices, with a cosine scheduler, starting with a learning rate of $1\times 10^{-5}$ and using 1000 warmup steps. The per-device batch size is set to 4 for both training and evaluation, and gradients are accumulated over 1 step. The maximum length of the output sequence is 2048 tokens. It takes 5 hours to train the model.
@@ A.3 Memory selection
In memory selection, the most appropriate memory must be chosen to ensure contextually relevant dialogue. To select the appropriate memory, it is necessary to consider all memories from previous sessions and all utterances from the current conversation. To prevent the model from being influenced by the order of the memories, and to mitigate any potential bias caused by sequence ordering, the memories are shuffled before being processed. The prompt used for memory selection is presented in Table 14. We use GPT-3.5-turbo for memory annotation and compare the performance of our trained memory selection model with GPT-3.5-turbo. As shown in Table 6, our model achieves a significantly higher F1 score, demonstrating the effectiveness of the learned memory selection model.
Model Recall Precision F1
GPT-3.5-Turbo 0.056 0.033 0.042
Ours 0.186 0.208 0.196
Table 6: Performance comparison between our trained memory selection model and GPT-3.5-turbo
@@ A.4 Information extraction
We conduct experiments to evaluate how similar and consistent the model’s output is compared to the actual ground truth. The experiments are performed using the BLEU-3/4 Papineni et al. (2002), ROUGE (ROUGE-1/2/L) Lin (2004), and BERTscore Zhang et al. (2019) metrics, and the model was compared against GPT-3.5-turbo. The results are presented in Table 7. The prompt used during training is very similar to the one used in the information extraction part of Session 3. However, in the training prompt, the phrase “from the movie {movie_name}. Use your knowledge of the characters to enhance your analysis where applicable.” is removed. The details of the modified prompt can be found in Table 11.
@@ A.6 Experiments
Evaluating the performance of open-domain conversations is critical for advancing natural language understanding. While traditional metrics remain important, recent methods have incorporated assessments utilizing models like GPT-4 to enhance evaluation accuracy. One such method, GPT Eval Liu et al. (2023), leverages Chain of Thought (CoT) reasoning to evaluate a model’s text generation. This technique offers greater context and structured guidance for large language models (LLMs) when assessing generated text. For this evaluation, we select multi-session dialogues from the SHARE dataset, each consisting of more than five sessions. From each session, two utterances are extracted to create dialogues within an EPISODE. The experiment is conducted five times for each criterion, and the results are averaged to determine the final outcome. The GPT Eval prompt used in this process is provide in Table 18, Table 19, Table 20, Table 21, Table 22, Table 23, Table 24.
@@ A.7 Human Evaluation
We conduct human evaluations through Amazon Mechanical Turk(AMT) to highlight the differences between SHARE and SHARE (w/o shared memory). We show the interface for the evaluation in Figure 4. Using a random seed of 2025, we randomly select 25 conversations from 96 test sets, which are constructed using the EPISODE Framework. We select annotators who are native English speakers and have a HIT Approval Rate of over 80% for all Requesters’ HITs. Then, we ask 20 annotators to evaluate each conversation based on the following criteria:
  • Closeness : In this criterion, we ask annotators to assess whether a response reflects the speakers’ familiarity, shared history, or understanding, expressed through teasing, emotional support, empathy, or references to past events, even during conflicts.
  • Engagingness : we ask the annotators to judge whether a response reveals emotional depth, relational tension, or moments of personal significance, even in calm exchanges.
  • Coherence : This criterion examines whether each response naturally connects to the preceding statements, even as the conversation evolves in different directions.
The evaluation takes place in the third session to maintain consistency in the testing environment. Table 8 provides a comparison between SHARE and SHARE (w/o shared memory) conversations, showing that SHARE achieves superior results over SHARE (w/o shared memory). This demonstrates that conversations with shared memory are more engaging and result in more natural dialogue.
@@ 4 Deduplicate information
Figure 5: The number of movies in each genre category in SHARE